OpenStudy (anonymous):

markus lawson is 6 yrs older than three times pam lincoln's age.the sum of their ages is 71.let a represent pam's age.

OpenStudy (anonymous):

markus lawson's age=m pam lincoln's age=a m=6+3a m+a=71 so: (6+3a)+a=71 6+4a=71 4a=65 a=16.25 m=54.75

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob